If a and b are two odd positive integers such that a>b, then how do we prove that one of the two numbers (a+b)/2 and (a-b)/2 is odd and the other is even?

Grade:11

1 Answers

Vikas TU
10 years ago

Take a = 5 and b = 3 such that a>b

Now, (a+b)/2 = (5+3)/2 = 4 which is even

and (a-b)/2 = (5-3)/2 = 1 which is odd

hence proved!

 

 

plz approve!